OpenStudy (anonymous):

What are the diameter and radius of a circle with a circumference of 48π yd? A. d = 24 yd r = 12 yd B. d = 24 yd r = 48 yd C. d = 12 yd r = 24 yd D. d = 48 yd r = 24 yd

OpenStudy (anonymous):

really need help

OpenStudy (anonymous):

were you going to type something

OpenStudy (jdoe0001):

\(\bf {\color{red}{ \textit{circumference of a circle}}}=2\pi r \\ \quad \\ {\color{red}{ 48\pi}}=2\pi r\implies \cfrac{48\pi}{2\pi}=r\qquad diameter=r+r\implies 2r \)

OpenStudy (anonymous):

it was D. d = 48 yd r = 24 yd got confused but thanks